Sutton Pointers Corporation expects to begin operations on January 1, 2015; it will operate as a specialty sales company that sells laser pointers over the Internet. Sutton expects sales in January 2015 to total $350,000 and to increase 15 percent per month in February and March. All sales are on account. Sutton expects to collect 66 percent of accounts receivable in the month of sale, 24 percent in the month following the sale, and 10 percent in the second month following the sale.

Required a. Prepare a sales budget for the first quarter of 2015.

b. Determine the amount of sales revenue Sutton will report on the first 2015 quarterly pro forma income statement.

c. Prepare a cash receipts schedule for the first quarter of 2015. (Do not round intermediate calculations. Round your answers to the nearest whole dollar amount.)

d. Determine the amount of accounts receivable as of March 31, 2015. (Do not round intermediate calculations. Round your answers to the nearest whole dollar amount.)
Answer #1

Answer-a)-

Sutton Pointers Corporation
Sales Budget
For the First Quarter of 2015
Particulars January February March
Sales $350000 $350000*1.15 =$402500 $402500*1.15 =$462875
Total Sales $350000 $402500 $462875

b)- Amount of sales revenue on the first 2015 quarterly pro forma income statement = $350000+$402500+$462875

= $1215375

c)-

Sutton Pointers Corporation
Schedule of Collections from sales
For the First Quarter of 2015
Particulars January February March
Total Sales $350000 $402500 $462875
Cash Collection
66% in month of sale $350000*66% = $231000 $402500*66%= $265650 $462875*66%= 305497.50
24% in following month of sale $350000*24% = $84000 $402500*24%= $96600
10% in the second month following sale $350000*10% = $35000
Total cash collection $231000 $349650 $437098

d)- The amount of accounts receivable as of March 31, 2015 =$197628.

Explanation- Amount of accounts receivable as of March 31, 2015 = February month accounts receivable+ March month accounts receivable

= ($402500*10%)+ ($462875*34%)

= $40250+157377.5

= $197627.50